# Foundation Models for the Brain and Body

### v2

@@ 2.1 Problem Definition
2.2 Unimodal Encoder Pre-training
We adopt the standard Transformer encoder vaswani2023attentionneed backbone with a linear tokenizer to map raw signals into $d$-dimensional token embeddings. A patch length of $S$ along the temporal dimension, results in $K=\left\lfloor\frac{T}{S}\right\rfloor$ non-overlapping tokens. We flatten the channels at each time-point within the patch and use linear tokenizer to project each patch into an embedding $\mathbf{z_{i}}\in\mathbb{R}^{d}$. Following Masked autoencoder (MAE) he2021maskedautoencodersscalablevision , given a mask ratio $r\in(0,1)$, we sample a random mask set $\mathcal{M}\subset\{1,\dots,K\},|\mathcal{M}|=K\times r$. The encoder $\phi$ processes only the unmasked tokens $\{\mathbf{z}_{i}:i\notin\mathcal{M}\}$, and the decoder $\psi$ reconstructs the full sequence. We optimize the masked-token reconstruction via mean squared error: $\mathcal{L}_{\mathrm{MAE}}=\;\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\bigl{\|}\psi\bigl{(}\phi(\{\mathbf{z}_{j}\}_{j\notin\mathcal{M}})\bigr{)}_{i}\;-\;\mathbf{z}_{i}\bigr{\|}_{2}^{2}.$ We apply this pre-training independently to the EMG and pose data, yielding two pre-trained encoders $\mathcal{E}_{x}=\phi_{\text{EMG}}$ and $\mathcal{E}_{p}=\phi_{\text{pose}}$, which produce meaningful representations for contrastive alignment.
@@ 2.3 Contrastive Pose-EMG Pre-training (CPEP)
In our approach, the hand pose data, offering high-quality, low-noise gesture information, serves as the anchor modality. In contrast, EMG signals exhibit lower signal-to-noise ratio (SNR). To align EMG and pose representations, we append a lightweight projection head $h_{\phi}$ to the pre-trained EMG encoder $\mathcal{E}_{x}$, mapping its [CLS] token into the same $d$-dimensional space as the pose encoder’s [CLS] token. The [CLS] token is chosen for its ability to capture global context and provide a compact summary of the entire sequence he2021maskedautoencodersscalablevision . During contrastive training, we keep the pose encoder $\mathcal{E}_{p}$ frozen and only update the EMG encoder $\mathcal{E}_{x}$ (together with the projection head $h$) to avoid the pose representations to be altered when enforcing alignment with EMG representations. For each paired sample $(x_{i},p_{i})$ we compute $u_{i}=h\!\left(\mathcal{E}_{x}(x_{i})\right)_{\mathrm{[CLS]}}$, $v_{i}=\left(\mathcal{E}_{p}(p_{i})\right)_{\mathrm{[CLS]}}$, where the projection head $h$ is randomly initialized and trained, and the pose encoder $\mathcal{E}_{p}$ is frozen. We $\ell_{2}$-normalize embeddings and use cosine similarity. Let $\tilde{u}_{i}=\frac{u_{i}}{\|u_{i}\|},\quad\tilde{v}_{j}=\frac{v_{j}}{\|v_{j}\|},\quad s_{ij}=\frac{\tilde{u}_{i}^{\top}\tilde{v}_{j}}{\tau},$ with temperature $\tau$. Following CLIP radford2021learningtransferablevisualmodels , the symmetric Information Noise-Contrastive Estimation (InfoNCE) objective oord2019representationlearningcontrastivepredictive is:
@@
This loss pulls EMG embedding $\tilde{u}_{i}$ toward its matching pose $\tilde{v}_{i}$ while pushing away all non-matching poses, yielding pose-informed EMG embeddings that support generalizable gesture classification.
@@ 2.4 Evaluation Protocols
We employ two evaluation protocols to examine the learned representation quality. Linear probing (LP): with labeled EMG data $\{(x_{i},y_{i})\}$, we freeze $\mathcal{E}_{x}^{*}$ and train a randomly-initialized linear classifier $\mathcal{C}$ on top of its embeddings, reporting accuracy on a held-out split. Zero-shot classification (ZS) proceeds analogously to CLIP’s image–text retrieval radford2021learningtransferablevisualmodels : for each EMG sample, we retrieve its TopK nearest poses in the embedding space, then vote the corresponding gesture labels to determine the predicted gesture. Given a test EMG sample $x_{j}$, let $\mathcal{R}_{j}=\operatorname*{TopK}_{p\in\mathcal{P}_{\mathrm{test}}}\!\bigl{(}\mathcal{E}_{x}(x_{j})^{\top}\mathcal{E}_{p}(p))\bigr{)}$ be the set of $k$ pose samples with highest cosine similarity to $x_{j}$. We then predict $\hat{y}_{j}=\mathop{\mathrm{mode}}\bigl{\{}\,y(p)\mid p\in\mathcal{R}_{j}\bigr{\}},$ where $y(p)$ denotes the gesture class associated with pose $p$.
@@ 3.1 Data
Dataset: We use emg2pose salter2024emg2poselargediversebenchmark , a large-scale open-source EMG dataset containing 370 hours of sEMG and synchronized hand pose data across 193 consenting users, 29 different behavioral groups that include a diverse range of discrete and continuous hand motions such as making a fist or counting to five. The hand pose labels are generated using a high-resolution motion capture system. The full dataset contains over 80 million pose labels and is of similar scale to the largest computer vision equivalents. Each user completed four recording sessions per gesture category, each with a different EMG-band placement. Each session lasted 45–120 s, during which users repeatedly performed a mix of 3–5 similar gestures or unconstrained freeform movements. We use non-overlapping 2-second windows as input sequences. EMG is instance-normalized, band-pass filtered (2–250 Hz), and notch-filtered at 60 Hz. For more details, please refer to salter2024emg2poselargediversebenchmark .
@@
Experiment Design: We evaluate on two disjoint gesture sets drawn from the public emg2pose corpus. First, we select four representative single-hand motions covering various finger movements as our in-distribution gestures. Second, from the six held-out classes that are not seen during training, we exclude the two-handed gesture and the highly variable “finger freeform” class, yielding four unseen gestures. Note that we include gestures like "CountingWiggling" and "FingerPinches" that are known to be challenging for vision-based hand pose estimation due to visual occlusion salter2024emg2poselargediversebenchmark . Details of gesture classes are in Appendix A. For data splits, we follow the public train $\mathcal{D}_{tr}$, val $\mathcal{D}_{val}$ , test $\mathcal{D}_{test}$ splits and define our data splits for downstream gesture classification tasks as shown in Table 1. The model is pre-trained on the full $\mathcal{D}_{tr}$. A linear head is trained on $\mathcal{D}_{\text{probe-tr}}^{\text{in}}$, and report final accuracy on the evaluation set $\mathcal{D}_{\text{eval}}^{\text{in}}$. For unseen gestures (which appear only in the original val and test splits), we train on $\mathcal{D}_{\text{probe-tr}}^{\text{unseen}}$, and report accuracy on $\mathcal{D}_{\text{eval}}^{\text{unseen}}$. Zero-shot classification is evaluated only on $\mathcal{D}_{\text{eval}}$. All users in $\mathcal{D}_{\text{eval}}$ are unseen, so the results also assess user-level generalization. A held-out dataset $\mathcal{D}_{\text{tune}}$, strictly disjoint from all other sets, is reserved for hyper-parameter tuning.
@@ 3.2 Gesture Classification Results
We assess the quality of EMG representations on both in-distribution and unseen gestures using the two evaluation protocols described earlier: zero-shot classification (ZS) and linear probing (LP). For comparative analysis, we compare our CPEP model against: (i) a supervised encoder–decoder Transformer called PoseT trained to regress poses directly from EMG; (ii) the supervised pose regression models (emg2pose, Vemg2pose, and NeuroPose) released with the emg2pose benchmark salter2024emg2poselargediversebenchmark ; and (iii) a self-supervised MAE model pre-trained using EMG data only (EMG-MAE). For a fair comparison under linear probing, we freeze each baseline’s encoder and train a softmax linear head on top of its embeddings. For context, we include an upper-bound by linear probing Pose-MAE features, since the pose encoder serves as the alignment anchor. Figure 2 presents a summary of the experimental results. We report macro accuracy (mean per-class accuracy) to account for class imbalance across gestures. We omit standard deviations over random seeds because they were negligible relative to the performance gains. CPEP outperforms all baselines in terms of accuracy under both linear probing and zero-shot settings. Remarkably, zero-shot classification on unseen gestures achieves performance comparable to linear probing, without requiring any additional task-specific training. Furthermore, our zero-shot CPEP approach exceeds the linear probing results of the baselines, highlighting its superior generalizability through contrastive learning. These findings support our claim that aligning EMG signals with high-quality pose embeddings enables effective generalization to unseen gestures.
@@ 3.3 Ablation Study
Impact of CPEP design choices. Table 2 summarizes the ablation results for different CPEP design choices. Initializing the EMG encoder from the pre-trained MAE consistently outperforms random initialization (CPEP-$\mathcal{E}_{x}$ RandInit) and results in a more stable contrastive training; random init slows convergence and degrades both zero-shot and linear-probing performance. One important finding is that randomly initializing both encoders ($\mathcal{E}_{x}$ and $\mathcal{E}_{p}$) fails to converge. Freezing both encoders and training only the projection head (CPEP-$\mathcal{E}_{x}$ Frozen) leads to significantly lower accuracy, highlighting the necessity of fine-tuning the EMG encoder during the feature alignment phase. Making the pose encoder trainable alongside the EMG encoder (updating both) also fails to converge, suggesting that altering the pose anchor harms alignment. Furthermore, average pooling over tokens (CPEP-AvgPool) underperforms the [CLS] token.
@@
Impact of $S_{\text{emg}}$, $S_{\text{pose}}$, and mask ratio $r$. We study the effect of the unimodal MAE pre-training choices (patch length $S$ and mask ratio $r$) on downstream classification, with CPEP training setup fixed. Fig. 3 summarizes the results across different hyper-parameter settings. Overall, for small values of $S_{\text{pose}}$, higher $r$ values achieves better results. With $S_{\text{pose}}{=}200$, r=0.5r{=}0.5 outperforming others. Using longer EMG patches consistently degrades performance.
@@ Appendix A Details of Gestures
(1) Finger pinches; (2) Thumb rotation; (3) Thumb swipes (whole hand); (4) Hand claw, grasp, and flicks.
@@
(1) Shaka and Vulcan peace; (2) Hook, OK, and Scissors; (3) Counting up/down; (4) Counting up/down with finger wiggling.
@@ Appendix B Implementation Details
We use 2 s windows sampled at 2 kHz for both pose and EMG. EMG is instance-normalized, band-pass filtered (2–250 Hz), and notch-filtered at 60 Hz. Following [16], we apply channel-rotation augmentation to EMG. Our MAE is a encoder–decoder Transformer model with 4 encoder layers and 2 decoder layers, and the embedding dimension is d=256d{=}256. We optimize with AdamW [11] (lr $1\mathrm{e}{-4}$, weight decay $1\mathrm{e}{-5}$) and cosine annealing with warm restarts [10]. The masking ratio is 50%. Token length is S=200S{=}200 for pose and S=50S{=}50 for EMG, producing non-overlapping tokens along time. Batch size is 256, and each MAE is trained for 100 epochs. Our supervised baseline PoseT share the same transformer architecture with MAE model, the only difference is that the outputs are pose regression predictions. The training objective is also the same as MAE, a mean squared error loss.

### v3

@@ 2.1 Problem Definition
2.2 Unimodal Encoder Pre‐training
We adopt the standard Transformer encoder vaswani2023attentionneed backbone with a linear tokenizer to map raw signals into dd‐dimensional token embeddings. A patch length of $S$ along the temporal dimension results in $K=\left\lfloor\frac{T}{S}\right\rfloor$ non‐overlapping tokens. We flatten the channels at each time-point within the patch and use linear tokenizer to project each patch into an embedding $\mathbf{z_{i}}\in\mathbb{R}^{d}$. Following Masked autoencoder (MAE) he2021maskedautoencodersscalablevision, given a mask ratio $r\in(0,1)$, we sample a random mask set $\mathcal{M}\subset\{1,\dots,K\},|\mathcal{M}|=K\times r$. The encoder $\phi$ processes only the unmasked tokens $\{\mathbf{z}_{i}:i\notin\mathcal{M}\}$, and the decoder $\psi$ reconstructs the full sequence. We optimize the masked‐token reconstruction via mean squared error: $\mathcal{L}_{\mathrm{MAE}}=\;\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\bigl\|\psi\bigl(\phi(\{\mathbf{z}_{j}\}_{j\notin\mathcal{M}})\bigr)_{i}\;-\;\mathbf{z}_{i}\bigr\|_{2}^{2}.$ We apply this pre-training independently to the EMG and pose data, yielding two pre-trained encoders $\mathcal{E}_{x}=\phi_{\text{EMG}}$ and $\mathcal{E}_{p}=\phi_{\text{pose}}$, which produce meaningful representations for contrastive alignment.
@@ 2.3 Contrastive Pose-EMG Pre-training (CPEP)
In our approach, the hand pose data, offering high-quality, low-noise gesture information, serves as the anchor modality. In contrast, EMG signals exhibit lower signal-to-noise ratio (SNR). To align EMG and pose representations, we append a lightweight projection head $h_{\phi}$ to the pre-trained EMG encoder $\mathcal{E}_{x}$, mapping its [CLS] token into the same $d$-dimensional space as the pose encoder’s [CLS] token. The [CLS] token is chosen for its ability to capture global context and provide a compact summary of the entire sequence he2021maskedautoencodersscalablevision. During contrastive training, we keep the pose encoder $\mathcal{E}_{p}$ frozen and only update the EMG encoder $\mathcal{E}_{x}$ (together with the projection head $h$) to avoid the pose representations being altered when enforcing alignment with EMG representations. For each paired sample $(x_{i},p_{i})$ we compute $u_{i}=h\!\left(\mathcal{E}_{x}(x_{i})\right)_{\mathrm{[CLS]}}$, $v_{i}=\left(\mathcal{E}_{p}(p_{i})\right)_{\mathrm{[CLS]}}$, where the projection head $h$ is randomly initialized and trained, and the pose encoder $\mathcal{E}_{p}$ is frozen. We $\ell_{2}$-normalize embeddings and use cosine similarity. Let $\tilde{u}_{i}=\frac{u_{i}}{\|u_{i}\|},\quad\tilde{v}_{j}=\frac{v_{j}}{\|v_{j}\|},\quad s_{ij}=\frac{\tilde{u}_{i}^{\top}\tilde{v}_{j}}{\tau},$ with temperature $\tau$. Following CLIP radford2021learningtransferablevisualmodels, the symmetric Information Noise-Contrastive Estimation (InfoNCE) objective oord2019representationlearningcontrastivepredictive is:
@@
This loss pulls EMG embedding $\tilde{u}_{i}$ toward its matching pose $\tilde{v}_{i}$ while pushing away all non-matching poses, yielding pose-informed EMG embeddings that support generalizable gesture classification. The embedding dimension is $256$ and CPEP was trained with a batch size of $256$.
@@ 2.4 Evaluation Protocols
We employ two evaluation protocols to examine the learned representation quality. Linear probing (LP): with labeled EMG data $\{(x_{i},y_{i})\}$, we freeze $\mathcal{E}_{x}^{*}$ and train a randomly-initialized linear classifier $\mathcal{C}$ on top of its embeddings, reporting accuracy on a held-out split. Zero‑shot classification (ZS) is performed as k-nearest-neighbor voting in the embedding space, following standard practice in representation learning (marks2024closerlookbenchmarkingselfsupervised; radford2021learningtransferablevisualmodels). For each EMG sample, we retrieve its TopK nearest poses in the embedding space, then vote the corresponding gesture labels to determine the predicted gesture. Given a test EMG sample $x_{j}$, let $\mathcal{R}_{j}=\operatorname*{TopK}_{p\in\mathcal{P}_{\mathrm{test}}}\!\bigl(\mathcal{E}_{x}(x_{j})^{\top}\mathcal{E}_{p}(p))\bigr)$ be the set of $k$ pose samples with highest cosine similarity to $x_{j}$. We then predict $\hat{y}_{j}=\mathop{\mathrm{mode}}\bigl\{\,y(p)\mid p\in\mathcal{R}_{j}\bigr\},$ where $y(p)$ denotes the gesture class associated with pose $p$.
@@ 3.1 Data
Dataset: We use emg2pose salter2024emg2poselargediversebenchmark, a large-scale open-source EMG dataset containing 370 hours of sEMG and synchronized hand pose data across 193 consenting users, 29 different behavioral groups that include a diverse range of discrete and continuous hand motions such as making a fist or counting to five. The hand pose labels are generated using a high-resolution motion capture system. The full dataset contains over 80 million pose labels and is of similar scale to the largest computer vision equivalents. Each user completed four recording sessions per gesture category, each with different EMG band placement. Each session lasted 45–120 s, during which users repeatedly performed a mix of 3–5 similar gestures or unconstrained freeform movements. We use non-overlapping 2-second windows as input sequences. EMG is instance-normalized, band-pass filtered (2–250 Hz), and notch-filtered at 60 Hz. For more details, please refer to salter2024emg2poselargediversebenchmark.
@@
Experiment Design: We evaluate on two disjoint gesture sets drawn from the public emg2pose corpus. First, we select four representative single‑hand motions covering various finger movements as our in‑distribution (in-dist.) gestures. Second, from the six held‑out classes that are not seen during training, we exclude the two‑handed gesture and the highly variable “finger freeform” class, yielding four unseen gestures. Note that we include gestures like “CountingWiggling” and “FingerPinches” that are known to be challenging for vision-based hand pose estimation due to visual occlusion salter2024emg2poselargediversebenchmark. Details of gesture classes are in Appendix A. For data splits, we follow the public train $\mathcal{D}_{tr}$, val $\mathcal{D}_{val}$ , test $\mathcal{D}_{test}$ splits in salter2024emg2poselargediversebenchmark and define our data splits for downstream gesture classification tasks as shown in Table 1. The model is pre-trained on the full $\mathcal{D}_{tr}$. A linear head is trained on $\mathcal{D}_{\text{probe-tr}}^{\text{in}}$, and report final accuracy on the evaluation set $\mathcal{D}_{\text{eval}}^{\text{in}}$. For unseen gestures (which appear only in the original val and test splits), we train on $\mathcal{D}_{\text{probe-tr}}^{\text{unseen}}$, and report accuracy on $\mathcal{D}_{\text{eval}}^{\text{unseen}}$. Zero-shot classification is evaluated only on $\mathcal{D}_{\text{eval}}$. All users in $\mathcal{D}_{\text{eval}}$ are unseen, so the results also assess user-level generalization. A held-out dataset $\mathcal{D}_{\text{tune}}$, strictly disjoint from all other sets, is reserved for hyper-parameter tuning.
@@ 3.2 Gesture Classification Results
We assess the quality of EMG representations on both in-dist. and unseen gestures using the two evaluation protocols described earlier: zero-shot classification (ZS) and linear probing (LP). For comparative analysis, we compare our CPEP model against: (i) a supervised encoder–decoder Transformer called PoseT trained to regress poses directly from EMG; (ii) the supervised pose regression models (emg2pose, Vemg2pose, and NeuroPose) released with the emg2pose benchmark salter2024emg2poselargediversebenchmark; and (iii) a self-supervised MAE model pre-trained using EMG data only (EMG-MAE). For a fair comparison under linear probing, we freeze each baseline’s encoder and train a softmax linear head on top of its embeddings. For context, we include an upper-bound by linear probing Pose-MAE features, since the pose encoder serves as the alignment anchor.
Figure 2 presents a summary of the experimental results. Since supervised baselines do not support zero-shot (k-nearest neighbor) classification in the embedding space, we report only their LP results. We report macro accuracy (mean per-class accuracy) to account for class imbalance across gestures. We omit standard deviations over random seeds because they were negligible relative to the performance gains. CPEP outperforms all baselines in terms of accuracy under both linear probing and zero-shot settings. Remarkably, zero-shot classification on unseen gestures achieves performance comparable to linear probing, without requiring any additional task-specific training. Furthermore, our zero-shot CPEP approach exceeds the linear probing results of the baselines, highlighting its superior generalizability through contrastive learning. These findings support our claim that aligning EMG signals with high-quality pose embeddings enables effective generalization to unseen gestures.
@@ 3.3 Ablation Study
Impact of CPEP design choices. Table 2 summarizes the ablation results for different CPEP design choices. Initializing the EMG encoder from the pre-trained MAE consistently outperforms random initialization (CPEP-$\mathcal{E}_{x}$ RandInit) and results in a more stable contrastive training. Random initialization slows convergence and degrades both zero-shot and linear-probing performance. One important finding is that randomly initializing both encoders ($\mathcal{E}_{x}$ and $\mathcal{E}_{p}$) fails to converge. Freezing both encoders and training only the projection head (CPEP-$\mathcal{E}_{x}$ Frozen) leads to significantly lower accuracy, highlighting the necessity of fine-tuning the EMG encoder during the feature alignment phase. Making the pose encoder trainable alongside the EMG encoder (updating both) also fails to converge, suggesting that altering the pose anchor harms alignment. Furthermore, average pooling over tokens (CPEP-AvgPool) underperforms the [CLS] token.
@@
Impact of $S_{\text{emg}}$, $S_{\text{pose}}$, and mask ratio $r$. We study the effect of the unimodal MAE pre-training choices (patch length $S$ and mask ratio $r$) on downstream classification, with CPEP training setup fixed. Fig. 3 summarizes the results across different hyper-parameter settings. Overall, for small values of $S_{\text{pose}}$, higher $r$ values achieve better results. With $S_{\text{pose}}{=}200$, r=0.5r{=}0.5 outperforming others. Using longer EMG patches consistently degrades performance.
@@ Appendix A Details of Gestures
(1) Thumb swipes whole hand; (2) Hand claw, grasp, and flicks; (3) Thumbs rotation, thumbs up and down; (4) finger pinches, single/multiple fingers.
@@
(1) Hook ’em Horns, OK, and Scissors; (2) Shaka and Vulcan peace; (3) Counting up/down; (4) Counting up/down with finger wiggling and spreading.
@@ Appendix B Implementation Details
We use 2 s windows sampled at 2 kHz for both pose and EMG. EMG is instance-normalized, band-pass filtered (2–250 Hz), and notch-filtered at 60 Hz. Following [salter2024emg2poselargediversebenchmark], we apply channel-rotation augmentation to EMG. Our MAE is an encoder–decoder Transformer model with 4 encoder layers and 2 decoder layers, and the embedding dimension is d=256d{=}256. We optimize with AdamW [loshchilov2019decoupledweightdecayregularization] (lr $1\mathrm{e}{-4}$, weight decay $1\mathrm{e}{-5}$) and cosine annealing with warm restarts [loshchilov2017sgdrstochasticgradientdescent]. The masking ratio is 50%. Token length is S=200S{=}200 for pose and S=50S{=}50 for EMG, producing non-overlapping tokens along time. Batch size is 256, and each MAE is trained for 100 epochs. Our supervised baseline PoseT shares the same transformer architecture with MAE model, the only difference is that the outputs are pose regression predictions. The training objective is also the same as MAE, a mean squared error loss.